OpenStudy (anonymous):

Square Root 2y-1-1=2 What is the root?

OpenStudy (anonymous):

The answers are Y=5 Y=4 or None

OpenStudy (anonymous):

Do you mean \(\sqrt{2y-1}-1=2\)?

OpenStudy (anonymous):

i do

OpenStudy (ddcamp):

First isolate the radical (part under the square root) by adding 1 to both sides: \[\sqrt{2y-1} = 3\]

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

I do? may I be the man of honor?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

heheh

OpenStudy (ddcamp):

Then, square both sides: \[2y-1 = 9\] And solve for y.

OpenStudy (anonymous):

so y = 5?